# Generalized Animal Imitator: Agile Locomotion with Versatile Motion Prior

### v2

@@ IV-A Evaluation of Learned Motion Priors
VIM $\mathbf{1.24\scriptstyle{\pm 0.62}}$ $0.01\scriptstyle{\pm 0.02}$ $0.11\scriptstyle{\pm 0.06}$ $\mathbf{0.08\scriptstyle{\pm 0.06}}$ $\mathbf{0.03\scriptstyle{\pm 0.03}}$ $\mathbf{13.313\scriptstyle{\pm 11.48}}$ $\mathbf{166.783\scriptstyle{\pm 120.217}}$
VIM (w/o Style Reward) $1.49\scriptstyle{\pm 0.69}$ $\mathbf{0.00\scriptstyle{\pm 0.01}}$ $0.12\scriptstyle{\pm 0.06}$ $5.14\scriptstyle{\pm 1.71}$ $0.25\scriptstyle{\pm 0.06}$ $6.28\scriptstyle{\pm 6.73}$ $109.76\scriptstyle{\pm 110.67}$
IV-B Evaluation on High-level Tasks
To further evaluate how our method leverages learned agile locomotion skills for High-level tasks, we designed a set of tasks and tested our method against standard baselines in both simulation and real-world scenarios.
High-level Tasks: Our task suite comprises: Following Command: This involves directing the robot to move with specific linear and angular velocities. The linear velocity command and angular velocity command are sampled uniformly between $0\sim 2$ m/s and $-2\sim 2$ rads/s respectively. In our motion prior, the robot is trained to move and turn at the reference motion’s speed; hence, to follow a command precisely, the high-level policy should smoothly interpolate between different speeds. Jump Forward: This task requires the robot to jump while running forward. We have adapted a subset of jumping rewards from CAJun [58] to evaluate policy interpolation between jumping and running motions within a fixed timeframe. Following Command + Jump Forward: Here, the robot must either jump forward or adjust to changing commanded speeds. To optimize episode return, the robot should not only use the agile locomotion skills from the reference motion dataset but also develop unobserved skills like executing sharp turns.
@@ IV-B Evaluation on High-level Tasks
Baselines: Considering the baseline’s subpar performance in low-level motion prior training, we compare our system with three representative baselines without pre-trained low-level controllers: PPO [45]: Demonstrates controllers trained exclusively on High-level task rewards. AMP [41] utilize existing reference motion to provide styling reward in an adversarial imitation learning manner and learn the policy for the High-level task while mimicking the behavior of reference motions. Hierarchical Reinforcement Learning (HRL) adapts from Jain et al. [25] which learns a high-level policy sending latent commands to a low-level motor controller. HRL resembles a broad category of prior works that decompose temporally extended reasoning into sub-problems [50, 11, 38, 35, 18, 28, 21]. For a fair comparison, we made modifications like removing the trajectory generator in [25], using PPO for AMP and HRL, and supplying full reference motion data to AMP and HRL integrated with AMP.
Evaluation in Simulation & Real World: We train all methods on each High-level task for $4\times 10^{8}$ environment samples with 3 random seeds. The simulation results are detailed in Figure 8, and real-world results are provided in Table IV. Real-world Following Commands trajectory is also provided in Figure 9. For the Following Command task, while all methods mastered basic locomotion, ours excelled in efficiency and smoothly transitioned between diverse linear and angular velocities. The other tasks, Jump Forward and Following Command + Jump Forward, demanded advanced jumping abilities, which baselines couldn’t emerge. These baseline methods either continuously moved forward, remained grounded when prompted to jump, or toppled to evade energy consumption penalties. In contrast, our system seamlessly bridged jumping and running actions, securing the highest episode return. Despite providing with a comprehensive reference motion dataset, baselines couldn’t harness the skills. This shortcoming possibly stems from the challenge of deriving agile locomotion skills from the dataset using only adversarial stylization rewards, mirroring the GAIL baseline’s poor performance in low-level motion prior training.
IV-C Ablation Study of Learned Motion Prior
To further understand how different reward and reference motions influence our framework, we provide further ablation study over reward terms as well as a more fine-grained performance of our framework when it imitates different reference motions.
When we remove the Functionality (Func) reward or Style reward, we found the performance of our method drops to a different extent, as shown in Table V . We found that without Funtionality reward, the learned controller could not robustly track the reference motion resulting in smaller Episode Return and, shorter Episode Length. On the other hand, removing Style reward results in a significantly higher Joint Angle and End-Effector tracking error.
In our framework, when the root pose of the simulated robot diverges too much from the reference root pose, we terminate the episode, as described in Sec III-E . In this case, the episode length is a good indicator of whether the learned policy could follow the reference motions. As shown in Fig. 10 , the performance of our framework varies when imitating different reference motions. When imitating relatively steady motions like Walk (Mocap), Pace (Mocap), Left Turn (Mocap), the learned controller could track the motion for a longer period. When imitating relatively agile motions, especially with high moving speed, such as Canter (Mocap), Jump While Running (Mocap), the performance of our system drops. This phenomenon is rooted in the methodology disparities between our robot and real animals.
V Conclusion
In this paper, we propose Versatile Instructable Motion prior (VIM) which learns agile locomotion skills from diverse reference motions with a single motion prior. Our results in simulation and in the real world show that our VIM captures both the functionality and the style of locomotion skills from reference motions. Our VIM also provides a temporally consistent and compact latent skill space representing different locomotion skills for different High-level tasks. With agile locomotion skills in our VIM, complex High-level tasks can be solved efficiently with minimum human effort.
Supplementary Material
I Detailed Observation Space
@@
  • Joint Angle - $\mathbb{R}^{12\times 3}$ contains joint rotations for all joints ($12$D) for the past three control step.
  • Joint Velocity - $\mathbb{R}^{12\times 3}$ contains joint velocities for all joints ($12$D) for the past three control step.
  • Previous Action - $\mathbb{R}^{12\times 3}$ contains positional command for all joints ($12$D) for the past three control step.
  • Projected Gravity - $\mathbb{R}^{3\times 3}$ contains the projected gravity in the robot frame, representing the orientation of the robot for the past three control steps.
  • Foot position - $\mathbb{R}^{3\times 4\times 3}$ contains the robot foot positions in the robot frame, $3$ dim per foot per timestep for the past three control steps
@@
We also provide additional high-level observation for the example high-level tasks we used. For Following Command task, we provide target linear velocity and target angular velocity as high-level observation. For Jumping Forward task, since the robot is tasked to jump forward in a fixed frequency, we provide normalized temporal phase as high-level observation.
II Domain Randomization

### v3

@@ Appendix E Single Skill Comparison in Simulation
Jump Forward (Optimization)
VIM $1.16\scriptstyle{\pm 0.52}$ $0.01\scriptstyle{\pm 0.01}$ $0.10\scriptstyle{\pm 0.05}$ $0.99\scriptstyle{\pm 0.30}$ $0.04\scriptstyle{\pm 0.01}$ $38.47\scriptstyle{\pm 9.18}$ $422.49\scriptstyle{\pm 96.00}$
Motion Imitation $1.19\scriptstyle{\pm 0.45}$ $0.00\scriptstyle{\pm 0.00}$ $0.13\scriptstyle{\pm 0.06}$ $4.24\scriptstyle{\pm 1.30}$ $0.12\scriptstyle{\pm 0.02}$ $26.11\scriptstyle{\pm 8.76}$ $325.08\scriptstyle{\pm 105.22}$
GAIL (Single Skill AMP) $2.00\scriptstyle{\pm 0.48}$ $0.04\scriptstyle{\pm 0.01}$ $0.10\scriptstyle{\pm 0.05}$ $0.92\scriptstyle{\pm 0.22}$ $0.03\scriptstyle{\pm 0.01}$ $11.71\scriptstyle{\pm 6.59}$ $161.38\scriptstyle{\pm 83.92}$
Jump While Running (Mocap)
VIM $1.58\scriptstyle{\pm 0.56}$ $0.01\scriptstyle{\pm 0.01}$ $0.09\scriptstyle{\pm 0.03}$ $1.63\scriptstyle{\pm 0.18}$ $0.06\scriptstyle{\pm 0.01}$ $13.36\scriptstyle{\pm 7.16}$ $172.90\scriptstyle{\pm 90.81}$
Motion Imitation $1.50\scriptstyle{\pm 0.49}$ $0.00\scriptstyle{\pm 0.00}$ $0.09\scriptstyle{\pm 0.03}$ $3.04\scriptstyle{\pm 0.99}$ $0.13\scriptstyle{\pm 0.06}$ $10.93\scriptstyle{\pm 5.08}$ $163.36\scriptstyle{\pm 74.99}$
GAIL (Single Skill AMP) $2.19\scriptstyle{\pm 0.84}$ $0.04\scriptstyle{\pm 0.01}$ $0.17\scriptstyle{\pm 0.04}$ $2.48\scriptstyle{\pm 0.61}$ $0.11\scriptstyle{\pm 0.02}$ $4.61\scriptstyle{\pm 2.84}$ $120.48\scriptstyle{\pm 81.14}$
Trot (Mocap)
VIM $1.21\scriptstyle{\pm 0.32}$ $0.00\scriptstyle{\pm 0.00}$ $0.08\scriptstyle{\pm 0.04}$ $0.18\scriptstyle{\pm 0.03}$ $0.01\scriptstyle{\pm 0.00}$ $21.52\scriptstyle{\pm 10.58}$ $213.10\scriptstyle{\pm 103.26}$
Motion Imitation $1.21\scriptstyle{\pm 0.29}$ $0.00\scriptstyle{\pm 0.00}$ $0.10\scriptstyle{\pm 0.05}$ $0.17\scriptstyle{\pm 0.03}$ $0.01\scriptstyle{\pm 0.00}$ $17.62\scriptstyle{\pm 8.86}$ $174.88\scriptstyle{\pm 87.40}$
GAIL (Single Skill AMP) $1.76\scriptstyle{\pm 0.78}$ $0.00\scriptstyle{\pm 0.00}$ $0.08\scriptstyle{\pm 0.05}$ $0.61\scriptstyle{\pm 0.60}$ $0.03\scriptstyle{\pm 0.03}$ $14.87\scriptstyle{\pm 8.76}$ $159.49\scriptstyle{\pm 90.19}$
Left Turn (Mocap)
VIM $0.07\scriptstyle{\pm 0.08}$ $0.00\scriptstyle{\pm 0.00}$ $0.16\scriptstyle{\pm 0.07}$ $0.15\scriptstyle{\pm 0.02}$ $0.01\scriptstyle{\pm 0.00}$ $31.64\scriptstyle{\pm 14.21}$ $299.63\scriptstyle{\pm 135.47}$
Motion Imitation $0.11\scriptstyle{\pm 0.12}$ $0.00\scriptstyle{\pm 0.00}$ $0.14\scriptstyle{\pm 0.07}$ $0.60\scriptstyle{\pm 0.41}$ $0.03\scriptstyle{\pm 0.02}$ $35.31\scriptstyle{\pm 14.05}$ $383.56\scriptstyle{\pm 135.10}$
GAIL (Single Skill AMP) $0.15\scriptstyle{\pm 0.16}$ $0.00\scriptstyle{\pm 0.00}$ $0.17\scriptstyle{\pm 0.08}$ $0.18\scriptstyle{\pm 0.10}$ $0.01\scriptstyle{\pm 0.01}$ $27.75\scriptstyle{\pm 16.92}$ $268.36\scriptstyle{\pm 162.53}$
Appendix F Implementation Details
Observation & Action Space: Our low-level observation includes joint angles, joint velocities, gravity vector in the robot frame, and the previously executed actions. Our controller outputs target joint angles for 12 joints of our robot in 25hz. The target joint angles are converted to torque command with PD controller where KP=40, KD=1.0.
@@ Appendix F Implementation Details
Reference Encoder $E_{ref}$ Proprioception Encoder $E_{prop}$: Our reference encoder proprioception encoder are both two-layer MLP with $[256]$ hidden units, mapping the reference motion segment into a $64$ dimensional latent distribution and proprioception into a $64$-dim robot state feature, respectively.
Low-level Policy $\pi_{low}$ and Value Network $V_{low}$: Our low-level policy is a three-layer MLP with $[256,128]$ hidden units, mapping the robot state feature and latent command to $12$-dim robot target joint angles. Our low-level value network shares the same structure while taking a motion embedding as additional input, and output 1-dim value for RL training. Our learnable motion embedding is a $64$-dim vector for each reference motion.
High-level Policy $\pi_{high}$ and Value Network $V_{high}$: Our high-level is formulated as a three-layer MLP with $[256,128]$ hidden units, mapping proprioception information and high-level task information to high-level latent command for low-level motion prior. Our high-level value network shares the same structure. High-level task information depends on specific task. Additional implementation details are provided in the supplementary materials
Simulation Setup: We utilize IsaacGym[68] to simulate 4096 robots in parallel and our simulation runs in 200 Hz. During motion prior training, for each robot, we uniformly sample a reference motion from the dataset for it to track.
Appendix G High Level Task Reward
Our high level jumpping reward is adapted from CAJun [59] with the following terms.
$r_{\text{jump}}$ $=2*(2-\left\lVert v_{\text{robot}}-2\right\rVert)/2+5*(\text{BaseHeight}-0.6)*\max(\sum_{f\in\mathrm{feet}}\hat{c}_{f}-4,0)$
$+3*\sum_{f\in\mathrm{feet}}\left\lVert 1+c_{f}-\hat{c}_{f}\right\rVert^{2}+1*\sum_{f\in\mathrm{feet}}\left\lVert c_{f}-\hat{c}_{f}\right\rVert*\min(h^{f},0.16)/0.16$
Here our desired foot contact $\hat{c}_{f}$ at each step is a binary value generated by the task generator as in CAJun [59] with value 0 for no contact, and value 1 for contact, similar for the actual contact $c_{f}$, and $h^{f}$ is the foot height over the ground.
Our high level command following reward is defined as follows.
$r_{\text{following cmd}}$ $=1.5*\exp\left(\left\lVert v_{\text{command}}-v_{\text{robot}}\right\rVert^{2}/0.25\right)+1.5*\exp\left(\left\lVert\omega_{\text{command}}-\omega_{\text{robot}}\right\rVert^{2}/0.25\right)$
$-2\left\lVert\text{Base Height}-\ \text{Target Height}\right\rVert$
Here the $v_{\text{command}}$ is the commanded target linear forward velocity, $v_{\text{robot}}$ is the current forward velocity of the robot. The $\omega_{\text{command}}$ is the commanded target angular velocity, $\omega_{\text{robot}}$ is the current angular velocity of the robot.
We also include energy penalty and action smooth reward as shown in the other rewards in Appendix F.
Appendix H Additional Low-level Skill Comparison
In addition to the low-level skill comparison in Figure 5, we provide another low-level skill comparison in Figure 9. Our controller learned to jump forward in the air with a natural gait, while baselines failed to leave the ground or failed to move forward
Appendix I High-level Policy Visualization
To better understand of the performance of our high-level policy, we provide Following Command trajectory in Figure 10. Though our low-level controller only learns to turn with specific angular velocity, our high-level could track different angular velocity in the real world.
Appendix J Detailed Observation Space
@@ Appendix I High-level Policy Visualization
$\bullet$ Joint Angle - $\mathbb{R}^{12\times 3}$ contains joint rotations for all joints ($12$D) for the past three control step.
$\bullet$ Joint Velocity - $\mathbb{R}^{12\times 3}$ contains joint velocities for all joints ($12$D) for the past three control step.
$\bullet$ Previous Action - $\mathbb{R}^{12\times 3}$ contains positional command for all joints ($12$D) for the past three control step.
$\bullet$ Projected Gravity - $\mathbb{R}^{3\times 3}$ contains the projected gravity in the robot frame, representing the orientation of the robot for the past three control steps.
$\bullet$ Foot position - $\mathbb{R}^{3\times 4\times 3}$ contains the robot foot positions in the robot frame, $3$ dim per foot per timestep for the past three control steps
@@
We also provide additional high-level observations for the example high-level tasks we used. For Following Command task, we provide target linear velocity and target angular velocity as high-level observations. For Jumping Forward task, since the robot is tasked to jump forward in a fixed frequency, we provide normalized temporal phase as high-level observation.
Appendix K Domain Randomization
@@
Appendix M Additional Results Regarding Stylization Tracking
Though adversarial training is generally unstable, we found it relatively stable during our training. We provide the training log of our discriminator and the average adversarial reward across epochs in Figure 11. Specifically, We applied the following techniques to stabilize the adversarial training.
  • We clipped the gradient of the discriminator to have the maximum norm of $0.5$
  • We applied gradient penalties during the training of the discriminator, following AMP [22]
We think our motion imitation reward also helped stabilize the adversarial training since our joint tracking and end-effector position tracking reward provide fine-grained instruction for policy training. We didn’t observe specific latent skill space collapse during our training, we think this is for the following reasons:
  • We provide motion embedding for the value function to distinguish different reference motions during training, as shown in Figure 3 in the manuscript. With motion embedding, the value function in our method could learn to distinguish different skills easily.
  • Our tracking reward terms provided dense instruction for the controller to generate different behaviors, which further regularized the latent skill space during training.
Figure 11: Adversarial Training Log
To study the training behavior in more detail, we visualized the episode return for the joint angle tracking term with the average adversarial reward in Figure 12. We found that in the first 1000 epoch, the episode return of joint tracking increased swiftly corresponding to the rapidly decreasing period of average adversarial reward. After 1000 epochs, the increasing rate of the episode return of joint tracking drops. We think this phenomenon corresponds to our claim in the manuscript that the model transits from learning the overall motion, where the episode return of joint tracking boost, towards learning the fine-grained behavior using the joint angle tracking reward, where the episode return of joint tracking grows slowly.
Figure 12: Average Adversarial Reward with Episode Return (Joint Angle Tracking)